OpenStudy (anonymous):

Suppose Mary is in consumer equilibrium. The marginal utility of good A is 30, and the price of good A is $2. a. if the price of good B is $4, the price of good C is $3, the price of good D is $1, and the price of all other goods and services is $5, what is the marginal utility of each goods mary is purchasing? b. If mary has chosen to keep $10 in savings, what is the ratio of MU to P for savings?

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

the answer involves simple mathematical equation which i may not be able to state explicitly because i don't know how to use the equation option. The requirement for equilibrium for two or more commodities is MUx/Px= MUy/Py .following this equilibrium condition to solve for all unknown MU of B, C, D and all other goods, lets say E. We have muA/PA=muB/pB=muC/pC=muD/pD=muE/pE. solving one by one,we have \[30/2=B/4\] \[B=30\times4/2\] B=120/2 muB= 60 YOU CAN GET OTHER VALUES OF muC,muD,muE by equating the already known values with them. u should get muC=45, muD=15, muE=75 .....BEST OF LUCKCOLLEAGUE